Ann is twice as old as Ben and thrice as old as Carrie. By the time Carrie is twice her age now, she will be half as old as Ann. If Ben is three years older than Carrie, how old is each one now? After how many years will be Carrie half as old as Ann?
Answer by lwsshak3(11628)

Ann is twice as old as Ben and thrice as old as Carrie. By the time Carrie is twice her age now, she will be half as old as Ann. If Ben is three years older than Carrie, how old is each one now? After how many years will be Carrie half as old as Ann?
***
let a=Ann's age
let b=Ben's age
let c=Carrie's age
..
a=3c
a=2b
b=c+3
..
a=2b=2(c+3)=2c+6
a=3c
3c=2c+6
c=6
..
a=3c=18
b=c+3=9
..
how old is each one now?
Carrie=6
Ben=9
Ann=18
..
After 6 years:(twice Carrie's age)
Carrie's age=6+6=12
Ann's age=18+6=24
After how many years will be Carrie half as old as Ann? 6